Presentation is loading. Please wait.

Presentation is loading. Please wait.

Status Report TOTEM collaboration meetings 24-MAY-2010

Similar presentations


Presentation on theme: "Status Report TOTEM collaboration meetings 24-MAY-2010"— Presentation transcript:

1

2 Status Report TOTEM collaboration meetings 24-MAY-2010

3 Overview DCS is in operation 24/7 DCS team activities Maintenance Development of new features

4 Generic Panel (improvements)

5 RP motorization panel (new panel) Shows motorization status for all the pots Log messages and configuration and special commands

6 RP pot panel (redesign) Shows only information of an specific pot Motorization and Silicon Detectors

7 LHC handshake (redesign) DIP and LHC Handshake are now isolated processes This should improve stability of the LHC Handshake Handshake problems reported to the JCOP support, and some things will be improved, but not everything

8 Self-Monitoring box ELMB boxes (under testing) Sensors box for T1

9 USB-ethernet adapters ELMB -> CANbus -> SYSTEC -> USB -> DIGI -> ethernet Will help in maintenance, if a computer crashes, all software can be redeployed to another node without having to recable in the counting room Already in production in the lab Will be used also in H8

10 Moving of the laboratory to 506 Everything ready and operational But took several working days…. of several people….

11 LHC information (done) Excel file for all the information we need (similar to pinout tables) Every week something is changing!!!! Integration of BPM ‘within days’

12 Consolidation activities (done) Exactly the same software in IP5 that in the laboratory As a ‘shortcut’ to provide something working, the computer names were hardcoded last year, so it was not possible to debug properly the system before sending to production. Migration of SVN repositories From bigbrother.cern.ch to the central services Fixes in the User Interface

13 DCS in H8 for RpSi Replication of the production software form IP5

14 Workforce confirmed 50%PPCERN 100%FLR3 yearsCERN 100%IA2.5 yearsCERN 20%FR1 yearCERN 50%JS0.5 yearBudapest 100%VT3 monthsTrainee Kuopio-Helsinki 100%JT3 monthsTrainee Kuopio 100%GP3 monthsSummer Student

15 Issues (from Jira) Last 30 days, not in absolute scale All maintenance effort on Ivan and Fernando Is good having trainees, but they need supervision They stay ‘short’ time have to focus in very specific tasks 13 new issues 7 issues resolved 98 issues open new/resolved issues date

16 Next steps (same list of Feb) Fix internal DCS problems (some parts have to be rewritten to add new functionality) Include in the pinout tables the LV max HW currents Finish LHC information integration RadMon readout Fix User Interface problems (to allow FSM operation) Propagation of states in the FSM tree (but no automatic actions) OK ?

17

18 DCS status summary (Feb 2010)

19 DCS status summary (May 2010)

20 Milestone Plan 2 WP closed 3 WP started -3 months Too many things on going

21 Activity Schedule D8 (done)

22 Activity Schedule D9 (in progress)

23 Activity Schedule I7 (in progress)

24 ?

25 THE END


Download ppt "Status Report TOTEM collaboration meetings 24-MAY-2010"

Similar presentations


Ads by Google